OpenStudy (anonymous):

help plzzz medal to whoever help me Which best describes the set of even numbers between one and eleven? A. {–10, –8, –6, –4, –2, 0, 2, 4, 6, 8, 10} B. {2, 4, 6, 8, 10} C. all real numbers D. null set

OpenStudy (anonymous):

help me plzzz

OpenStudy (anonymous):

B! What do you need help with?

OpenStudy (anonymous):

5 questions

OpenStudy (anonymous):

Okay, I'll help! What are the rest? First though, want me to explain this one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so "even" numbers are basically those that can be divided by 2, and still return an integer [counting number, but also includes negatives, like 1, 2, 3, 4 and -1, -2, -3, -4, etc.]. Now, in set notation, we have a set that describes all of the numbers that are even [divisible by two] and lead from 1 to 11. First, here's the set of all numbers from 1 to 11: n[I chose to assign "n", but choose whatever you want to]={1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11}. Okay, now here's the set of even numbers [going to infinite essentially]: e[once again, I chose e, but do what you want]={2, 4, 6, 8, 10, 12, 14, 16, 18, 20, ... , etc.}. Now, we want the Union of these 2 sets [essentially, the set that contains all of the numbers that are in BOTH sets; here is how you write that: nUe [Union of sets n and e] = {2, 4, 6, 8, 10}. Any questions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

heres the other onee

OpenStudy (anonymous):

Okay, so "even" numbers are basically those that can be divided by 2, and still return an integer [counting number, but also includes negatives, like 1, 2, 3, 4 and -1, -2, -3, -4, etc.]. Now, in set notation, we have a set that describes all of the numbers that are even [divisible by two] and lead from 1 to 11. First, here's the set of all numbers from 1 to 11: n[I chose to assign "n", but choose whatever you want to]={1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11}. Okay, now here's the set of even numbers [going to infinite essentially]: e[once again, I chose e, but do what you want]={2, 4, 6, 8, 10, 12, 14, 16, 18, 20, ... , etc.}. Now, we want the Union of these 2 sets [essentially, the set that contains all of the numbers that are in BOTH sets; here is how you write that: nUe [Union of sets n and e] = {2, 4, 6, 8, 10}. Any questions?

OpenStudy (anonymous):

What is the intersection of the sets A = {2, 3, 5, 7, 11} and B = {5, 11, 13, 15, 17}? A. {2, 3, 5, 7, 11, 13, 15, 17} B. {5, 11} C. {2, 3, 5, 7, 11} D. null set

OpenStudy (anonymous):

OKAY, I MADE A MISTAKE. When I said Union, I actually meant to write "intersection". An Intersection of 2 sets is the combination of elements in BOTH sets. I apologize. The Intersection of sets n and e [represented as an Upside Down U] = {2, 4, 6, 8, 10}. Is this clear before I answer your above question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the other is still right??

OpenStudy (anonymous):

Okay, so what we want for your above question, is the set that contains elements ONLY that are in BOTH sets. Do you think you can do it?

OpenStudy (anonymous):

Yes, the other is still correct. I just accidentally used the wrong term. I accidentally said "Union" but I meant to say "Intersection". Intersection is when you only include elements from BOTH sets.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

so now we on this question

OpenStudy (anonymous):

What is the intersection of the sets A = {2, 3, 5, 7, 11} and B = {5, 11, 13, 15, 17}? A. {2, 3, 5, 7, 11, 13, 15, 17} B. {5, 11} C. {2, 3, 5, 7, 11} D. null set

OpenStudy (anonymous):

@faysal96 y u leave

OpenStudy (anonymous):

@dan815 can u help me

OpenStudy (anonymous):

Okay, so since set A = {2, 3, 5, 7, 11}, and set B = {5, 11, 13, 15, 17}, take the elements that are only in BOTH, {5, 11}. Does this make sense to you?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So basically, a set is a bunch of numbers, right? Think of it like a bucket filled with different fruits [apples, oranges, pineapples, grapefruits, watermelons, etc.]. So we have 2 buckets, and let's say the first bucket has an apple, a grapefruit, a watermelon, a banana, an orange, and a pineapple. The second bucket has a watermelon, a kiwi, a banana, a pineapple, and a coconut. "What is the intersection of these 2 sets?" This is really just asking, what do both buckets have in common? The answer is a watermelon, a banana, and a pineapple! So the intersection of Bucket 1 {apple, grapefruit, watermelon, banana, orange, pineapple} and Bucket 2 {watermelon, kiwi, banana, pineapple, coconut}, is the following set: {watermelon, banana, pineapple}. Does this make sense? The above question is basically the same thing, but you are just using numbers instead of names of fruit!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im on another 1 now

OpenStudy (anonymous):

Okay, anything else?

OpenStudy (anonymous):

What is the intersection of the three sets: A = {0, 2, 3, 6, 8}, B = {2, 3, 6, 8, 9}, and C = {1, 2, 4, 8, 9}? A. {2, 8, 9} B. {2, 6, 8} C. {2, 8} D. {0, 1, 2, 3, 4, 6, 8, 9}

OpenStudy (anonymous):

So again, take what you see that is in ALL THREE sets, and make that a new set. Can you do it?

OpenStudy (anonymous):

Like what's in common with all 3 sets?

OpenStudy (anonymous):

This table shows the top athletes that play both football and baseball. What is the intersection of the set of football players and the set of baseball players? 02_02_LQ_5.gif A. Intersection: {Jackson, Peavey} B. Intersection: {Peavey} C. Intersection: {Jackson} D. Intersection: {Adams}

OpenStudy (anonymous):

Again, take a look at the left hand side [the set of all football players] and the right hand side [set of all baseball players] and find what's in common with BOTH. Does that make sense?

OpenStudy (anonymous):

yes so is it ABCD

OpenStudy (anonymous):

Jackson and Peavey, so that would be A.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ty i turned it in lets se wat i get

OpenStudy (anonymous):

i got 100%

OpenStudy (anonymous):

would u like to help me on some more

OpenStudy (anonymous):

Sure! What do you need help with?

OpenStudy (anonymous):

im looking

OpenStudy (anonymous):

comparing expressions

OpenStudy (anonymous):

5 questions

OpenStudy (anonymous):

would u help me??

OpenStudy (anonymous):

Which value for x makes the open sentence true? 8 + 3 • x = 22 + x2 A. 5 B. 4 C. 3 D. 2

OpenStudy (anonymous):

@faysal96 do u see it??

OpenStudy (anonymous):

Yeah. is that an \[x^{2}\] at the end?

OpenStudy (anonymous):

omg hold up wat u say

OpenStudy (anonymous):

is the last term an x-squared?

OpenStudy (anonymous):

its x2

OpenStudy (anonymous):

ill repost the top

OpenStudy (anonymous):

Which value for x makes the open sentence true? 8 + 3 • x = 22 + x2 A. 5 B. 4 C. 3 D. 2

OpenStudy (anonymous):

Ooh, like as in it's the tens place? or is it 2*x?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

u want me to repost it again

OpenStudy (anonymous):

Which value for x makes the open sentence true? 8 + 3 • x = 22 + x2

OpenStudy (anonymous):

A. 5 B. 4 C. 3 D. 2

OpenStudy (anonymous):

there now it should be better

OpenStudy (anonymous):

@faysal96 do u understand it better

OpenStudy (anonymous):

this one is 5 question and we r on 1

OpenStudy (anonymous):

I don't know how to do this :/

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which expression shows how 8 • 54 can be rewritten using the distributive property? A. 4 • 50 + 4 • 4 B. 8 • 50 + 8 • 4 C. 8 • 50 + 8 • 54 D. 50 • 8 + 50 • 4

OpenStudy (anonymous):

8 * 54 = 432 8*50+8*4=432

OpenStudy (anonymous):

Which are the like terms in 4x3 – 7x2 + x + x3? A. 4x3, –7x2, and x3 B. 4x3and x C. 4x3and x3 D. They are all like terms.

OpenStudy (anonymous):

Only x.

OpenStudy (anonymous):

Simplify: 4x3 – 3x2 + x + 3x3 A. 5y3 B. 4x3– 3x2+ x + 3x3 (the expression cannot be simplified) C. 7x6– 3x2+ x D. 7x3– 3x2+ x